# Undesirable Memorization in Large Language Models: A Survey

### v1

@@ VI Mitigating memorization: strategies and techniques
As discussed in earlier sections, memorization is influenced by a range of factors and it is sometimes necessary for the learning process [94]. However, in scenarios where memorization could lead to privacy concerns or security vulnerabilities, some methods could be employed to mitigate its impact. In these cases, specific strategies are utilized to limit the retention of sensitive information, ensuring that potential risks related to data exposure or misuse are minimized.
VI-A Data de-duplication
Lee et al. [62] run exact matching and approximate matching (MiniHash) de-duplication algorithms on the C4, RealNews, LM1B, and Wiki40B datasets and show that these datasets contain up to 13.6% near duplicates and up to 19.4% exact duplicates; showing that de-duplication can reduce training cost in the first place. Measuring the amount of duplicates shared between train and validation splits of these datasets, they show that there is a possibility of validation data leakage which causes the over-estimation of the model accuracy or may lead to biases when considering hyperparameters tuning.
@@ VI-A Data de-duplication
To investigate the impact of data de-duplication on language model’s memorization, they train a 1.5B parameter GPT-2 model from scratch on three different settings: C4-Original, C4-NearDup, and C4-ExactSubstr, each for two epochs. Then the evaluate the memorization in no prompt and prompted settings and count the 50-token substrings generations that are presented in the training data as memorization. The no-prompt experiment generations show 10×10$\times$10 × less memorization in de-duplicated trained models. On the other hand, in the prompted experiment, when the prompt comes from the duplicate examples, the model trained on C4-Original generates the true exact continuation over 40% of the time. The other two models also generate the ground truth more often when the prompt is sampled from the duplicate examples, suggesting that more harsh de-duplication algorithms are needed to prevent memorization.
VI-B MemFree decoding
Ippolito et al. [32] try to prevent memorization using the novel MemFree Decoding method. Their approach proposes a new sampling strategy to prevent the emission of memorized samples. In the sampling phase, they keep track of the current n-gram (where ‘n‘ is a fixed parameter and should be chosen carefully; small n would prevent many common phrases to appear and big n may not prevent memorization) that is being generated and after generation of each token, they check the existence of the last n-gram in the training set. If this n-gram is a member of the training set, they ignore the newly generated token and sample another token, so as to obtain a non-memorized string. They inspect the GPT-Neo [95] family of models from smallest size to the largest one, prompting with 50 token sequences from the training data and anticipating for a 50 token suffix match. They show that their method significantly reduces the similarity of the generations with respect to the training data. They also show that although the approximate memorization increases with the model size log-linearly, MemFree defence stays effective in keeping the similarity around 0.6 BLEU score even in the largest model size.
While this verbatim memorization filtering reduces the average similarity score between the training data and the generations, qualitative analysis of the generations shows that the models opt to cheat in some sense. Ippolito et al. [32] inspect the production model GitHub Copilot [96] and show that when it is prompted with the training data, it successfully avoids verbatim generation, however, a simple style transfer in the prompts easily bypasses the Copilots memorization defence. They repeat this experiment on the English language models such as PaLM [36] and show how this defense could be bypassed, concluding this defence is an incomplete defence.
VI-C Differential privacy
@@
Differential privacy (DP) is a technique that protects individual data points by adding noise, minimizing the impact of any single data point on the model’s output. DP-SGD (Differentially Private Stochastic Gradient Descent) is an adaptation of the standard SGD algorithm, designed to fine-tune language models while maintaining privacy [97]. Carlini et al. [30] demonstrate that by adjusting the privacy budget parameter $\epsilon$ in DP-SGD training, the exposure of memorized data can be reduced to a level that makes it indistinguishable from any other data. However, this comes at the cost of reduced model utility and a slower training process.
@@
While DP is effective for reducing memorization and preserving privacy, it is crucial to select hyperparameters wisely. Otherwise, the model may not withstand stronger privacy attacks, potentially compromising its effectiveness as shown in Lukas et al. [59].
VI-D Unlearning methods
As discussed in the intentional forgetting section [IV-H], unlearning methods offer effective strategies to mitigate memorization by selectively removing specific information from trained models without the need for full retraining. However, a limitation of these unlearning methods is that they require prior knowledge of what the model has memorized in order to selectively remove it, which can be challenging to determine and may limit the effectiveness of these techniques.
VII Future directions and open challenges
@@
VII-A Balancing performance with privacy in LLMs
@@
As discussed earlier, memorization in LLMs can also have beneficial uses, such as improving factual recall and enhancing performance in specific tasks. The precondition is that privacy leakage and copyright issues have been taken care of. Methods like Differential Privacy (DP), that are designed to mitigate memorization risks and prevent sensitive data exposure, often come at the cost of reduced model performance, as they limit the model’s ability to retain useful information. To fully harness the benefits of memorization while safeguarding against privacy breaches, future research needs to focus on strategies that balance these competing goals. This includes developing techniques that protect sensitive data and intellectual property without significantly degrading the model’s accuracy and utility. Such efforts will be key to ensuring both ethical compliance and high-performance outcomes in LLMs.
VII-B Reducing factual memorization in favor of conceptual memorization
@@
The interplay between factual and conceptual memorization is relevant in LLM development, because there is a trade-off between correctness of information, and undesirable recall of training data: While factual memorization can provide precise information recall, it also presents higher risks in terms of privacy and data protection. Conceptual memorization, on the other hand, contributes to the model’s ability to generalize and apply knowledge flexibly, but may be more challenging to audit and control. For example, when using LLMs for question answering, literal reproduction of facts is desired; this includes names and numbers (e.g. “Who was the first person to fly across the ocean and when did this take place?”). It is straightforward to evaluate LLMs for generating the correct facts. If we prefer conceptual memorization in LLMs, we should accept a higher rate of hallucinations, because the model more freely generates text based on its parametric knowledge.
@@
Future research in this area should therefore focus on developing techniques to balance these two types of memorization, enhancing the benefits of each while mitigating their respective risks. This could involve methods to selectively encourage conceptual memorization while limiting unnecessary factual memorization, particularly of sensitive information, depending on the type of task at hand.
VII-C The boundary of memorization and understanding
@@
It is relatively straightforward to design experiments that demonstrate a model’s ability to generate fluent and human-like text, such as generating novel content or adapting to new contexts. However, verifying that these capabilities are not simply the result of memorization rather than abstraction over the information processed during training is much more challenging, as both can produce similar outcomes. Distinguishing between these requires carefully designed experiments, and further research is needed to clarify when a model is abstracting over learned concepts versus merely memorizing and reproducing them.
VII-D Memorization in specific contexts
@@
LLMs that have been fine-tuned for conversations can be used as conversational agents (chatbots), e.g., to assist customers of online services [102, 103]. Nasr et al. [34] introduce their so-called “divergence attack” to extract memorized samples from conversation-aligned LLMs; however, as they also discuss, these attacks are not powerful enough to stimulate training data reproduction. This makes us conclude that attacking conversation-aligned language models requires more advanced methods. Since most of the production language models are only available in a conversational settings, addressing memorization in conversational agents and conducting novel attacking methods to alleviate training data extraction in these models would be a prominent research direction.
@@
In RAG frameworks, LLMs are helped with a retrieval component that first selects the most relevant documents from a collection and feeds them to the prompt [104, 105]. The LLM then generates an answer based on the provided sources. This has advantages such as the reduction of hallucination [106], the transparency of sources, and the potential of generating answers related to proprietary or novel sources that were not in the LLM training data. Although very popular, RAG is not yet thoroughly analyzed regarding memorization. This is important, because a reduction of hallucination when using RAG in LLMs could also have an increase in memorization as a side effect: less hallucination means that the LLM stays closer to the original content. Recently, Zeng et al. [93] analyzed privacy aspects in RAG. They show that RAG systems are vulnerable of leaking private information. On the other hand, they also found that RAG can mitigate the reproduction of the LLM training data. A broader study focusing on memorization in RAG-based LLMs that compares different proposed retrieval architectures is an important future research direction.
@@
xLMs are trained to interpret and generate text in multiple languages. Trained on extensive datasets that include various languages, these models develop language-agnostic representations [107]. Despite their potential, xLMs often face challenges related to data scarcity for low-resource languages, leading to performance disparities [108]. This necessitates ongoing research to enhance their efficacy and fairness across all languages. For future work, we propose to investigate whether a low-resource language setting is more prone to memorization by comparing to English scenarios. Addressing this research direction is essential not only for understanding the privacy implications associated with LLMs and data leakage in low-resource settings but also for ensuring AI safety in societies using LLMs with languages that have fewer resources.
@@
Since DLMs show remarkable performance in the vision domain, researchers have started to adopt the diffusion models (DM) idea to the text domain and utilize their generative capabilities [109]. Carlini et al. [110] have conducted data extraction analysis on the image diffusion models, showing that diffusion models are much less private than prior generative models such as GANs, and that mitigating these vulnerabilities may require new advances in privacy-preserving training. Gu et al. [111] also show that according to the training objective of the diffusion models, a memorization behavior is theoretically expected and then they quantify the impact of the influential factors on the memorization behaviors in DMs. However, research focusing on the memorization issues related to DLMs for text remains unexplored. Even though the idea of diffusion language models is the same as the vision domain, they are inherently different, because of the discrete nature of the text domain. Therefore an analysis of the general vision diffusion models may not be applicable to the diffusion language models, making an independent research on memorization against DLMs a prominent future direction.
VIII Conclusion
Figure 3: The distribution by publication year of the papers discussed in this survey
@@ VIII Conclusion
In this paper, we systematically organized, summarized, and discussed the existing scientific work related to unintended and undesirable memorization in LLMs. Undesirable memorization might lead to privacy risks and other ethical consequences. We found that there exists a large body of research on memorization in LLMs, given the young age of the technology: Transformer models were first developed in 2017 [112], and the first generative large language model with emergent abilities was released in 2022 [1, 5]. The majority of papers on the topic were therefore published in very recent years (see Figure 3). This indicates that the field is working fast to analyze memorization and developing methods to mitigate undesirable memorization.
@@
Despite the fast-growing body of literature on the topic, we argue that there are important areas that require more attention in research in the coming years. We have pointed out four specific contexts in which memorization needs to be studied and, when needed and possible, mitigated: LLM-based conversational agents, retrieval-augmented generation, multilingual LLMs, and diffusion language models. These areas all are of high importance, not only from the academic perspective, but maybe even more from the application and industry perspective. In particular conversational agents and retrieval-augmented generation are actively being developed in the commercial context. When applied to proprietary databases or in interaction with customer information, these applications are particularly vulnerable to privacy and security risks.
@@
The main limitation of a SoK paper on a highly active research topic is the fast pace in which the field is moving. A survey paper written in 2024 risks to become outdated by 2026. Although we acknowledge this, we argue that the topic of memorization is too important to not create this systemization of knowledge – especially because of the added value that we created with our concrete suggestions for future directions.

### v2

@@ V Mitigating memorization: strategies and techniques
As discussed in earlier sections, memorization is influenced by a range of factors and it is sometimes necessary for the learning process [111, 112]. However, in scenarios where memorization could lead to privacy concerns or security vulnerabilities, some methods could be employed to mitigate its impact. In these cases, specific strategies are utilized to limit the retention of sensitive information, ensuring that potential risks related to data exposure or misuse are minimized.
V-A Data de-duplication
Lee et al. [77] run exact matching and approximate matching (MiniHash) de-duplication algorithms on the C4 [113], RealNews [114], LM1B [115], and Wiki40B [116] datasets and show that these datasets contain up to 13.6% near duplicates and up to 19.4% exact duplicates.
@@ V-A Data de-duplication
To investigate the impact of data de-duplication on a language model’s memorization, they train a 1.5B parameter GPT-2 [117] model from scratch on three different settings: C4-Original, C4-NearDup, and C4-ExactSubstr, each for two epochs. Then they evaluate the memorization in no-prompt and prompted settings and measure the 50-token exact match (Section III-A). The no-prompt experiment generations show $10\times$ less memorization in de-duplicated trained models. On the other hand, in the prompted experiment, when the prompt comes from the duplicate examples, the model trained on C4-Original generates the true exact continuation over 40% of the time. The other two models also generate the ground truth more often when the prompt is sampled from the duplicate examples, suggesting that more harsh de-duplication algorithms are needed to prevent memorization.
V-B Differential privacy
@@
Differential privacy is a data privacy method that ensures the results of any analysis over a dataset reveals minimal information about any individual’s data, protecting against privacy breaches [118]. This method is adopted in some techniques in machine learning to protect individual data points by adding noise, minimizing the impact of any single data point on the model’s output. DP-SGD (Differentially Private Stochastic Gradient Descent) is an adaptation of the standard SGD algorithm, designed to fine-tune language models while maintaining privacy [119]. Carlini et al. [24] demonstrate that by adjusting the privacy budget parameter $\epsilon$ in DP-SGD training, the exposure of memorized data can be reduced to a level that makes it indistinguishable from any other data. However, this comes at the cost of reduced model utility and a slower training process.
@@
While adopting the DP method in data privacy is mathematically proven to protect individuals’ privacy, it is crucial to select hyperparameters wisely when used in training LLMs as a technique; otherwise, the model may not withstand stronger privacy attacks, potentially compromising its effectiveness as shown in Lukas et al. [69].
V-C Unlearning methods
As memorization could lead to privacy risks and copyright issues, unlearning methods could be necessary in some situations. Methods like knowledge unlearning aim to selectively remove specific information from trained models without retraining them from scratch. Bourtoule et al. [124] introduce the “SISA” framework for efficient machine unlearning, which divides the training data into shards (shards partition the data into disjoint segments) and trains sub-models that can be easily retrained if data needs to be removed. For LLMs specifically, Chen and Yang [125] introduce lightweight unlearning layers into transformers, allowing for selective data removal without full model retraining. Pawelczyk et al. [126] introduce “In-Context Unlearning,” which involves providing specific training instances with flipped labels and additional correctly labeled instances as inputs during inference, effectively removing the targeted information without updating model parameters. Kassem et al. [127] propose “DeMem,” a novel unlearning approach leveraging a reinforcement learning feedback loop with proximal policy optimization to reduce memorization. By fine-tuning the model with a negative similarity score as a reward signal, the approach encourages the LLM to paraphrase and unlearn pre-training data while maintaining performance.
Additionally, knowledge unlearning techniques have been categorized into parameter optimization, parameter merging, and in-context learning, each offering unique advantages in efficiently removing harmful or undesirable knowledge from LLMs [128]. These methods not only enhance privacy and security but also ensure that the overall performance of the models remains intact, making them scalable and practical for real-world applications [129].
Unlearning methods offer effective strategies to mitigate memorization by selectively removing specific information from trained models without the need for full retraining. However, these methods are generally designed to target specific pieces of information, which means they rely on the ability to identify what the model has memorized beforehand. This reliance poses a challenge, as determining the exact content a model has memorized can be difficult and may limit the applicability of unlearning techniques for general mitigation against memorization.
V-D Heuristic Methods
Liu et al. [130] propose the alternating teaching method, a teacher-student framework where multiple teachers trained on disjoint datasets supervise a student model in an alternating fashion to reduce unintended memorization. This approach demonstrates superior privacy-preserving results on the LibriSpeech dataset [131] while maintaining minimal utility loss when sufficient training data is available. Ippolito et al. [26] propose “MemFree Decoding,” a novel sampling strategy to reduce memorization during text generation by avoiding the emission of token sequences matching the training data. They demonstrate its effectiveness across multiple models, including GPT-Neo [132] and Copilot [133], significantly lowering generation similarity with the training data. However, they also reveal that a simple style transfer in prompts can bypass this defense, highlighting its limitations. Similarly, Borec et al. [107] investigate the impact of nucleus sampling on memorization, showing that while increasing the nucleus size slightly reduces memorization, it only provides modest protection. They also highlight the distinction between “hard” memorization, involving verbatim reproduction, and “soft” memorization measured by the ROUGE similarity metric [134], where generated outputs echo the training data without exact replication. Hans et al. [135] introduce the “goldfish loss,” a subtle modification to the next-token training objective where randomly sampled subsets of tokens are excluded from the loss computation. This prevents models from memorizing complete token sequences, significantly reducing extractable memorization while maintaining downstream performance.
Answer to RQ4. Data deduplication has been widely shown to be one of the most effective and efficient methods to mitigate memorization. Differential privacy, unlearning, and other heuristic approaches can also be used to prevent memorization but those need to be employed precautiously as they might hurt the model performance.
VI Future directions and open challenges
@@
VI-A Balancing performance and privacy in LLMs
@@
As discussed earlier, memorization in LLMs can also have beneficial uses, such as improving factual recall and enhancing performance in specific tasks such as question answering. Privacy and copyright issues are some of the major concerns here [19, 20, 11]. Privacy-enhancing technologies such as multi-party computing, homomorphic encryption, and differential privacy, which can be employed to mitigate memorization risks and prevent sensitive data exposure, often come at the cost of reduced model performance, as they limit the model’s ability to retain precise information or perturb the model’s output. To fully harness the benefits of memorization while safeguarding against privacy breaches, future research needs to focus on strategies that balance these competing goals. This includes developing techniques that protect sensitive data and intellectual property without significantly degrading the model’s accuracy and utility. Such efforts will be key to ensuring both legal compliance and high-performance outcomes in LLMs.
VI-A1 Memorization and Differential Privacy
As discussed in Section III-D, counterfactual memorization shares some conceptual similarities with differential privacy. Counterfactual memorization evaluates the difference in a model’s behavior on a specific data point when the data point is included in or excluded from the training dataset. In contrast, a model is considered differentially private for a given example if the outputs of two models trained on neighboring datasets—differing by only that example—are indistinguishable.
In the context of text data, the definition of a “data point” can vary significantly. It might refer to named entities, individual sentences, complete documents, or even the entire text corpora. Both of these points underscore the need for further research to examine the interplay and dynamics between memorization and differential privacy at different data granularities.
VI-B Reducing verbatim memorization in favor of content memorization
@@
The interplay between factual and content memorization is relevant in LLM development, because there is a trade-off between the correctness of information, and undesirable recall of training data: Although verbatim memorization can provide more precise information recall, it also presents higher risks in terms of privacy and data protection. Content memorization, on the other hand, contributes to the model’s ability to generalize and apply knowledge flexibly, but it may also present more challenges for audit and control. For example, when using LLMs for question answering, literal reproduction of facts is desired; this includes names and numbers (e.g., “Who was the first person to fly across the ocean and when did this take place?”). It is straightforward to evaluate LLMs to generate the (correct) facts. If content memorization is preferred over verbatim memorization in LLMs, a higher rate of hallucinations should be accepted because the model generates text more freely based on its parametric knowledge.
@@
Future research in this area should therefore focus on developing techniques to balance these two types of memorization, enhancing the benefits of each while mitigating their undesired consequences. This could involve methods to selectively encourage content memorization while limiting unnecessary factual memorization, particularly of sensitive information, depending on the type of task at hand.
VI-C The boundary between memorization and understanding
@@
It is relatively straightforward to design experiments that demonstrate a model’s ability to generate fluent and human-like text, such as generating novel content or adapting to new contexts. However, verifying that these capabilities are not simply the result of memorization rather than the abstraction/generalization over the information processed during training is much more challenging, as both can produce similar outcomes. Distinguishing between these requires carefully designed experiments, and further research is needed to clarify when a model makes abstractions over learned concepts versus merely memorizing and reproducing them.
VI-D Memorization in specific contexts
@@
LLMs that have been fine-tuned for conversations can be used as conversational agents (chatbots), e.g., to assist customers of online services [136, 137]. Nasr et al. [40] introduce their so-called “divergence attack” to extract memorized samples from conversation-aligned LLMs. As they also discuss, these attacks are not powerful enough to stimulate training data reproduction. However, this does not mean that the conversation-aligned LLMs are not vulnerable to attacks or manipulation. It also means This makes us conclude that attacking conversation-aligned language models requires more advanced methods. Since most of the production language models are only available in a conversational settings, addressing memorization in conversational agents and conducting novel attacking methods to alleviate training data extraction in these models would be a prominent research direction.
@@
In RAG frameworks, LLMs are helped with a retrieval component that first selects the most relevant documents from a collection and feeds them to the prompt [138, 139]. The LLM then generates an answer based on the provided sources. This has advantages such as the reduction of hallucination [140], the transparency of sources, and the potential of generating answers related to proprietary or novel sources that were not in the LLM training data. Although very popular, RAG is not yet thoroughly analyzed regarding memorization. This is important, because a reduction of hallucination when using RAG in LLMs could also have an increase in memorization as a side effect: less hallucination means that the LLM stays closer to the original content. Recently, Zeng et al. [141] analyzed privacy aspects in RAG. They show that RAG systems are vulnerable to leaking private information. On the other hand, they also found that RAG can mitigate the reproduction of the LLM training data. A broader study focusing on memorization in RAG-based LLMs that compares different proposed retrieval architectures is an important future research direction.
@@
xLMs are trained to interpret and generate text in multiple languages. Trained on extensive datasets that include various languages, these models develop language-agnostic representations [142]. Despite their potential, xLMs often face challenges related to data scarcity for low-resource languages, leading to performance disparities [143]. This necessitates ongoing research to enhance their efficacy and fairness across all languages. In line with the significant effect of training data on memorization (discussed in section IV-B), for future work, we propose to investigate whether a low-resource language setting is more prone to memorization by comparing it to English scenarios. Addressing this research direction is essential not only for understanding the privacy implications associated with LLMs and data leakage in low-resource settings but also for ensuring AI safety in societies using LLMs with languages that have fewer resources.
@@
Since DLMs show remarkable performance in the vision domain, researchers have started to adopt the diffusion models (DM) idea to the text domain and utilize their generative capabilities [144]. Carlini et al. [145] have conducted data extraction analysis on the image diffusion models, showing that diffusion models are much less private than prior generative models such as GANs, and that mitigating these vulnerabilities may require new advances in privacy-preserving training. Gu et al. [146] also show that according to the training objective of the diffusion models, a memorization behavior is theoretically expected and then they quantify the impact of the influential factors on the memorization behaviors in DMs. However, research focusing on the memorization issues related to DLMs for text remains unexplored. Even though the idea of diffusion language models is the same as the vision domain, they are inherently different, because of the discrete nature of the text domain. Therefore, an analysis of the general vision diffusion models may not be applicable to the diffusion language models, making an independent research on memorization against DLMs a prominent future direction.
Answer to RQ5. Key aspects of memorization in LLMs still require exploration: Balancing privacy with performance remains challenging, as does reducing verbatim recall while preserving factual accuracy. Distinguishing memorization from true understanding needs better evaluation methods. The interplay between memorization and differential privacy at different granularities is underexplored. Additionally, specific contexts like conversational agents, RAG systems, multilingual models, and diffusion language models require deeper study to understand their unique memorization risks.
VII Conclusion
@@
In this paper, we organized, summarized, and discussed the existing scientific work related to undesirable memorization in LLMs. Undesirable memorization might lead to privacy risks and other ethical consequences. We found that there exists a large body of research on memorization in LLMs, given the young age of the technology: Transformer models were first developed in 2017 [147], and the first generative large language model with emergent abilities was released in 2022 [1, 5]. The majority of papers on the topic were therefore published in recent years. This indicates that the field is working fast to analyze memorization and developing methods to mitigate undesirable memorization.
@@
Despite the fast-growing body of literature on the topic, we argue that there are important areas that require more attention in research in the coming years. We have pointed out four specific contexts in which memorization needs to be studied and, when needed and possible, mitigated: LLM-based conversational agents, retrieval-augmented generation, multilingual LLMs, and diffusion language models. These areas all are of high importance, not only from the academic perspective, but maybe even more from the application and industry perspective. In particular, conversational agents and retrieval-augmented generation are actively being developed in the commercial context. When applied to proprietary databases or in interaction with customer information, these applications are particularly vulnerable to privacy and security risks.
@@
The main limitation of a survey paper on a highly active research topic is the fast pace at which the field is evolving. A survey paper written in 2025 risks becoming outdated by 2026. Although we acknowledge this, we argue that the topic of memorization is too important to overlook, especially given the added value we provide through our concrete suggestions for future directions. To mitigate this issue, we maintain a dedicated GitHub repository [^10] https://github.com/alistvt/undesirable-llm-memorization that catalogs the references discussed in this survey and will be regularly updated to reflect the latest developments in the field.